# Robust fuzzy clustering with cellwise outliers

### v1

@@ Abstract
Fuzzy clustering is a technique for identifying subgroups in heterogeneous populations by quantifying unit membership degrees. The magnitude of the latter depends on the desired level of fuzzification, based on the purpose of the analysis. We combine the advantages of fuzzy clustering with a robust approach able to detecting cellwise outliers, i.e., anomalous cells in a data matrix. The proposed methodology is formulated within a probabilistic framework and estimated via an Expectation-Maximization algorithm for missing data. It includes an additional step for flagging contaminated cells, which are then treated as missing information. The strengths of the model are illustrated through two real-world applications: the first one identifies individuals at potential risk of obesity based on their physiological measurements, while the second one analyzes well-being across regions of the OECD countries. We also explore the effects of the model’s tuning parameters and provide guidance for users on how to set them suitably.
@@
Cellwise contamination , Fuzzy assignments , Constrained estimation , Robust clustering , EM algorithm
@@ 1 Introduction
The goal of the unsupervised clustering is to discover subpopulations, called clusters, within the data which share common characteristics, and to estimate their statistical features (e.g., centers, scatter matrices, etc.). Distance-based methods include $k$-means [1, 2] as example, while in the probabilistic framework, one of the most well-known approaches is that of finite mixture models [3]. These methodologies have been extended to the fuzzy clustering setting [4], in which observations are not fully assigned to a specific cluster, but can have a positive membership degree to more than one cluster (see [5] for fuzzy $c$-means, and [6, 7, 8] for maximum likelihood approaches). Unlike finite mixture models, which provide a measure of uncertainty through the estimation of posterior probabilities – typically used to compute the Maximum A Posteriori (MAP) and assign each observation to a single cluster (hard assignment) – fuzzy approaches allow for varying degrees of fuzzification depending on the purpose of the analysis. This flexibility is especially important in domains such as medical diagnosis, treatment planning for subgroups of patients, and remote sensing. A concrete example comes from medical studies on certain diseases, where some patients may present overlapping symptoms or ambiguous laboratory test results, making it difficult to assign them uniquely to a diagnostic category. In such cases, fuzzy clustering enables, on the one hand, the identification of a subset of patients for whom further testing may be particularly recommended, and on the other hand, the possibility of adjusting the size of this subset based on the purpose of the analysis or specific clinical priorities. However, as this example highlights, the interest in fuzzification typically concerns only a subset of observations rather than all of them: for many observations, the membership to a specific cluster may be sufficiently strong that a hard assignment is appropriate. This motivates the use of fuzzy clustering methods with high contrast [9], where hard and soft assignments coexist within the same framework.
[…]
@@
Until the first decade of the 2000s, the casewise paradigm was the dominant framework for identifying outliers. However, the increasing dimensionality of the data has made it reasonable to assume that units may exhibit only a few outlying cells, while the remaining ones still contain reliable information. This has led to the development of the cellwise contamination model [16], where the goal is to flag as individual anomalous cells of a data matrix, rather than entire observations. In this setting, even a single outlying value can contaminate an observation, and as the number of variables increases, cellwise contamination is likely to affect nearly all observations. This causes clustering methodologies such as TCLUST and F-TCLUST to break down due to the high contamination of the cases. Additionally, in the cellwise paradigm, it is possible to take advantage of the reliable information within an observation to correct unreliable cells through imputation, despite discarding them for inference. Following this rationale, Raymaekers and Rousseeuw introduced the cellwise MCD (cellMCD, [17]) for estimating the location and scale parameters in a single population. In cellMCD, the cells flagged as outlying values are treated as Missing At Random [18] and are imputed during the E-step of the Expectation-Maximization (EM) algorithm [19] used for its implementation, which fits naturally within the missing data framework. cellMCD was extended to heterogeneous populations through a Gaussian mixture model approach, resulting in cellGMM [20] for unsupervised clustering and cellgGMM [21] for a semi-supervised approach, in which information on the clustering structure is known a priori. Further cellwise methodologies have been proposed, based on either discarding or imputing approaches, with goals different from clustering and/or beyond the scope of this paper, such as dimensionality reduction (see [22] for an overview). However, none of these methods addresses cellwise contamination in the context of fuzzy clustering.
In this paper, we aim to fill this gap by combining the cellwise outlier detection with the advantages of fuzzy clustering. We propose a cellwise fuzzy clustering method, called cellFCLUST, which intends to uncover a clustering structure that accommodates both hard and soft assignments on the one hand, and to robustly estimate the cluster parameters in the presence of cellwise contamination on the other. The proposal is formulated in the maximum likelihood framework, assuming a Gaussian distribution for the clusters and allowing each to have a distinct shape. By setting a proportion of unreliable cells for the variables, cellFCLUST identifies and imputes them using the information contained in the remaining reliable cells of each observation, making the use of an EM-type algorithm suitable for its estimation. As a result of the imputation, all observations are assigned to a cluster – unlike trimming approaches, which discard units, or noise clustering methods, where outliers are not grouped with regular data. The potential of the proposed methodology is demonstrated through real data sets from two different fields. In the first application, we identify groups of individuals with similar circumference measurements, which support the classification of subjects according to different body fat levels. The fuzzy clustering approach enables herein the detection of individuals whose values lie near the cluster boundaries, thereby highlighting those who may be considered at risk (e.g., individuals in the overweight or obese range). The second application concerns regional data for studying well-being in the OECD countries. It provides insights into the heterogeneity of living conditions across countries and highlights regions that may deviate in certain domains from the expected well-being patterns of their countries.
@@
The rest of the paper is organized as follows. Section 2 introduces cellFCLUST, its parameter estimation, and the corresponding algorithm. In Section 3, we provide guidance on the tuning parameter selection through examples that illustrate their impact on cluster recovery and outlier detection. Two real data applications are presented in Section 4. Finally, Section 5 concludes the paper with a discussion on the proposed methodology, as well as potential directions for future developments in the cellwise clustering literature.
@@ 2 Cellwise fuzzy clustering
Let $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\prime}$ be a data matrix composed of $n$ units, where $\mathbf{x}_{i}$ is a random vector taking values in $\mathbb{R}^{J}$. Some cells $x_{ij}$ of $\mathbf{X}$ can be missing or contaminated, generally referred to as unreliable cells. We denote by $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{n}]^{\prime}$ the $(n\times J)$ cellwise indicator matrix, where $w_{ij}=1$ if the cell is reliable, i.e., neither missing nor outlying, and $w_{ij}=0$ otherwise. Differently from the outlying values, the missing values are known, and therefore the corresponding zeros into $\mathbf{W}$ can be set a priori. According to $\mathbf{w}_{i}$, we partition $\mathbf{x}_{i}$ into $\mathbf{x}_{i[\mathbf{w}_{i}]}$ and $\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}$, where $\mathbf{w}_{i}^{c}=\mathbf{1}_{J}-\mathbf{w}_{i}$ with $\mathbf{1}_{J}$ being the unitary vector. These two sub-vectors represent the reliable and unreliable cells for the $i$th observation, respectively.
Our goal is to group the $n$ units into $K$ clusters using a fuzzy approach while handling contaminated data - and missing values, if present - and imputing them based on reliable information. To track the membership of an observation to a cluster, we define the $(n\times K)$ matrix $\mathbf{U}$, whose elements $u_{ik}$ take values into $[0,1]$ and such that $\sum_{k=1}^{K}u_{ik}=1$ for $i=1,\ldots,n$. We define a fuzzy clustering problem with cellwise outlier detection, called cellwise Fuzzy Clustering (cellFCLUST), in a maximum likelihood framework by the maximization of the following objective function
@@
where $m$ is the fuzzifier tuning parameter, $\phi_{J}(\cdot)$ is the probability density function of a $J$-variate normal distribution, and $J[\mathbf{w}_{i}]$ is the number of reliable cells for the $i$th observation. Both $\boldsymbol{m}_{1},\ldots,\boldsymbol{m}_{K}$, which are vectors in $\mathbb{R}^{J}$, and $\boldsymbol{\mathbf{S}}_{1},\ldots,\boldsymbol{\mathbf{S}}_{K}$, which are positive definite matrices of order $J$, are restricted in (1) to the subvectors and the submatrices, respectively, corresponding to the reliable cells for the $i$th observation. The fuzzifier parameter $m$ must be greater than $1$ to obtain strictly fuzzy memberships, since $m=1$ results in crisp assignments for every observation, even though $u_{ik}\in[0,1]$, as shown in [9]. The maximization of (1) is subject to the following constraints
@@
where $h$ and $c\geq 1$ are fixed constant, and $\lambda_{1}(\boldsymbol{\mathbf{S}}_{k}),\ldots,\lambda_{J}(\boldsymbol{\mathbf{S}}_{k})$ are the $J$ eigenvalues of $\boldsymbol{\mathbf{S}}_{k}$. Constraint (2) imposes that $h$ cells per variable are reliable. This value is set to $\left\lceil(1-\alpha)n\right\rceil$, where $\alpha$ represents the contamination level and should be at most $0.25$ to guarantee that pairs of variables overlap for some units, allowing their covariances to be computed. The eigenvalue-ratio constraint in (3) prevents unboundedness and spurious solutions that may arise during the maximization of the objective function. As the constant $c$ - which constraints the ratio between the largest and the smallest eigenvalues of the covariance matrices across clusters - decreases, the cluster configurations become increasingly similar, reaching spherical clusters when $c=1$. cellFCLUST turns out not to be affine equivariant due to the eigenvalue-ratio constraint. However, increasing $c$ yields a methodology that is nearly affine equivariant, although this increases the risk of finding spurious solutions. Moreover, cellFCLUST is not equivariant under nonsingular linear or orthogonal transformations, as the cells are tied to the coordinate system; that is, these transformations alter the position of the outlying cells, thereby distorting their identification.
It is worth noting that (1) may favor clusters with comparable values of $\sum_{i=1}^{n}u_{ik}^{m}$, which correspond to cluster sizes in a hard clustering approach. To avoid this effect and allow the sizes to vary across clusters, we consider the maximization of the following objective function
@@
where $p_{k}\in(0,1)$, such that $\sum_{k=1}^{K}p_{k}=1$, are weights to estimate. In the remainder of the paper, we consider unbalanced clusters and therefore focus on the maximization of (4). In Section 3, we provide evidence on the difference in the clustering structure resulting from the maximization of (1) rather than (4), particularly when the clusters have different sizes.
@@ 2.1 An EM-type algorithm for cellFCLUST
We solve the maximization problem of the objective function in (4) via an algorithm composed of four alternating steps, which is similar to an EM algorithm for missing data in the clustering framework [23]. Its key feature lies in the first step, called Concentration step, where nhn-h cells per variable considered as contaminated are flagged and set to zero in $\mathbf{W}$. The membership values in $\mathbf{U}$ are computed in Step 2. Following the EM rationale and treating the contaminated cells as missing, the parameters of the distribution for the missing part, necessary for the overall parameter estimation, are obtained in Step 3, resulting in the imputation of the contaminated cells, rather than their discard. Finally, Step 4 estimates the vectors $\boldsymbol{m}_{k}$ and the matrices $\boldsymbol{\mathbf{S}}_{k},k=1,\ldots,K$, subject to constraint (3). A detailed description of the algorithm is provided below.
@@
    Initial solutions for the parameters are obtained using several applications of the TCLUST method. Specifically, TCLUST is applied to each variable and pairs of variables to initialize $\mathbf{W}$, and then on random subsets of variables to achieve feasible initializations for $p_{k},\boldsymbol{m}_{k}$, and $\boldsymbol{\mathbf{S}}_{k},k=1,\ldots,K$ (see [20] for more details). Accordingly, an initial solution for $\mathbf{U}$ is computed as described in Step 2.
@@
    Given the current parameters and the actual configuration of $\mathbf{W}$, we update the latter column-by-column. Let consider the contribution of the $i$th unit to the objective function in (4), i.e.
@@
    where $\hat{x}_{ij(k)}$ and $C_{ij(k)}$ are the expectation and the variance, respectively, of the cell corresponding to the $j$th variable in the $i$th unit, conditional on the other reliable cells for the $i$th unit and considering the parameters of the $k$th cluster. By sorting the $\Delta_{ij}$ values in an increasing order, i.e., $\Delta_{(1)j}\leq\Delta_{(2)j}\leq\cdots\leq\Delta_{(n)j}$, we flag as unreliable the units with indexes $\{i:\Delta_{ij}<\Delta_{(n-h)j}\}$, where $h=\left\lceil(1-\alpha)n\right\rceil$, and set the corresponding elements of the $j$th column of $\mathbf{W}$ to zero. Equivalently, the units with $\{i:\Delta_{ij}\geq\Delta_{(n-h)j}\}$ are considered reliable, which is denoted by ones into the $j$th columns of $\mathbf{W}$.
    If missing values occur for the $j$th variable, the corresponding cells of $\mathbf{W}$ are set to zero. Consequently, the $\Delta$ values and the threshold $h$ are computed only on the observed cells. This approach may result in varying proportions of zeros across variables, depending on the extent of missingness they contain. However, using $\alpha$ to account for both the contaminated and missing cells – and therefore grounding $h$ on $n$, independently of the number of missing values – could potentially lead to an incorrect identification of the truly contaminated cells. In general, we recommend excluding variables with a high proportion of missing data, particularly when moderate contamination is expected.
@@
    $\begin{cases}u_{ik}=I\{f_{ik}=\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}\}&\text{if}\,\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}\geq 1\\ u_{ik}=\Bigg{(}\sum_{k^{\prime}=1}^{K}\Bigg{(}\dfrac{\log(f_{ik})}{\log(f_{ik^{\prime}})}\Bigg{)}^{\frac{1}{m-1}}\Bigg{)}^{-1}&\text{if}\,\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}<1\end{cases},$ (7)
[…]
@@
    Thanks to the separability assumption of the parameters and the Gaussian distribution properties, we can decompose the probability density function of the complete data $\phi_{J}(\mathbf{x}_{i};\boldsymbol{m}_{k},\boldsymbol{\mathbf{S}}_{k})$ into the product $\phi_{J[\mathbf{w}_{i}]}(\mathbf{x}_{i[\mathbf{w}_{i}]};\boldsymbol{m}_{k[\mathbf{w}_{i}]},\boldsymbol{\mathbf{S}}_{k[\mathbf{w}_{i},\mathbf{w}_{i}]})\times\phi_{J[\mathbf{w}_{i}^{c}]}(\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}|\mathbf{x}_{i[\mathbf{w}_{i}]};\boldsymbol{m}_{k[\mathbf{w}_{i}^{c}|\mathbf{w}_{i}]},\boldsymbol{\mathbf{S}}_{k[\mathbf{w}_{i}^{c},\mathbf{w}_{i}^{c}|\mathbf{w}_{i}]})$. Given the cellwise indicator matrix and the current parameters, we compute the parameters of the distribution of $\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}$ conditional on $\mathbf{x}_{i[\mathbf{w}_{i}]}$ as
@@
    If the update of the covariance matrices in (12) does not satisfy the eigenvalue-ratio constraint in (3), we apply the efficient procedure proposed by Fritz et al. [24]. This ensures that the constraint is met.

### v2

@@ Abstract
In a data matrix, it can be distinguished between observations, each represented by a full row vector for an individual, and cells, which correspond to single entries of that matrix. Recent developments in robust statistics have introduced the cellwise contamination paradigm, which assumes contamination on cells rather than on entire observations. This approach becomes particularly relevant as the number of variables increases. Indeed, discarding or downweighting entire observations because of a few anomalous values in them, as done by traditional (casewise) robust methods, can result in substantial information loss, since the non-contaminated (or reliable) cells can still be highly informative. This philosophy can also be considered in fuzzy clustering, by assuming that reliable cells within an observation may still provide useful information for determining fuzzy memberships. A robust fuzzy clustering proposal is thus introduced in this work, combining the advantages of dealing with outlying cells and simultaneously controlling the degrees of fuzziness of observation assignments. The cluster-specific relationships among variables, detected by the fuzzy clustering approach, are also key to better identifying outlying cells. The strengths of the proposed methodology are illustrated through a simulation study and two real-world applications. The effects of the model’s tuning parameters are explored, and some guidance for users on how to set them suitably is provided.
@@
Cellwise contamination , Fuzzy assignments , Constrained estimation , Robust clustering , High contrast property
journal: International Journal of Approximate Reasoning
@@ 1 Introduction
The goal of the unsupervised clustering is to discover subpopulations, called clusters, within the data which share common characteristics, and to estimate their statistical features (e.g., centers, scatter/covariance matrices, etc.). In many cases, clusters are not perfectly separated, and the information provided by hard assignments of the observations may offer a narrow perspective on the underlying structure of the data. To overcome this limitation of hard clustering algorithms such as $k$-means [1, 2], fuzzy clustering approaches have been introduced [3], in which observations are not fully assigned to a single cluster but can have positive membership degrees to more than one cluster. Both distance-based methods, which assume a spherical shape for the clusters (e.g., fuzzy $c$-means [4]), and probabilistic methods, in which this restriction is relaxed (e.g., maximum likelihood approaches [5, 6, 7, 8]), are examples of fuzzy clustering procedures. One of their main advantages is that they allow for varying degrees of fuzzification through a tuning parameter, which can be set depending on the purpose of the analysis. This feature cannot be achieved by other clustering methodologies, such as finite mixture models [9], which, although they quantify membership uncertainty through the estimation of posterior probabilities, do not provide control over the degree of fuzziness.
@@
The flexibility of the fuzzy clustering methods is especially important in domains such as medical diagnosis, treatment planning for subgroups of patients, and remote sensing. A concrete example comes from clinical studies on certain diseases, where some patients may present overlapping symptoms or ambiguous laboratory test results, making it difficult to assign them uniquely to a diagnostic category. In such cases, fuzzy clustering enables, on the one hand, the identification of a subset of patients for whom further testing may be particularly recommended, and on the other hand, the possibility of adjusting the size of this subset based on the purpose of the analysis or specific clinical priorities. However, as this example highlights, the interest in fuzzification typically concerns only a subset of observations rather than all of them: for many observations, the membership to a specific cluster may be sufficiently strong that a hard assignment is appropriate. This motivates the use of fuzzy clustering methods with high contrast [10], where hard and soft assignments coexist within the same framework.
In real-world applications, such as those previously mentioned, data may contain missing values, as well as measurement errors, anomalies, or, more generally, outliers. Missing information usually affects cells, i.e., variable measurements for a given observation, of the data matrix, whose position is known a priori. Several methodologies have been extended to cope with incomplete datasets, such as fuzzy $c$-means [11] (see [12] for an overview). On the other hand, outliers were traditionally defined as entire observations that do not follow the pattern of the majority of the data, therefore referred to as casewise or rowwise outliers [13], which are not known a priori. For instance, in the fuzzy clustering literature, Fuzzy TCLUST (F-TCLUST, [14]) was introduced to estimate the location and scatter parameters of the clusters using a maximum likelihood approach, while detecting and removing anomalous observations. This is related to the Minimum Covariance Determinant (MCD) estimator introduced by Rousseeuw [15, 16] in the single-population framework. F-TCLUST relies on the individual contribution to the objective function to identify a subset of observations that are retained as reliable for parameter estimation. Other previous alternative approaches can be found in [17], where fuzzy $c$-means is robustified via impartial trimming, and in [18], in which outliers are not discarded but instead modeled as belonging to a “noise” cluster.
@@
Until the first decade of the 2000s, the casewise paradigm was the dominant approach for identifying outliers. However, the increasing dimensionality of the data has made it reasonable to assume that observations may exhibit only a few outlying cells, while the remaining ones still contain reliable information. This has led to the development of the cellwise contamination model [19]. In the latter, discarding or downweighting entire observations, as done by traditional casewise robust methods, can result in substantial information loss, even when the overall proportion of atypical cells is very small (see Figure 1). Additionally, as the number of variables increases, cellwise contamination is likely to affect nearly all observations, often through a single outlying value in many of them, causing clustering methodologies such as F-TCLUST and those based on noise clusters to break down. Moreover, in the cellwise paradigm, it is possible to take advantage of the reliable information within an observation to obtain a predicted value for each cell and compare it with the observed one. If the former deviates from the latter, the cells are flagged as contaminated, and corrected through imputation for the parameter estimation, despite discarding them. Following this rationale, Raymaekers and Rousseeuw introduced the cellwise MCD (cellMCD, [20]) for estimating the location and scatter parameters in a single population, by treating the cells flagged as outlying values as Missing At Random (MAR, [21]). cellMCD has been extended in several contexts (see, for instance, [22], [23], and [24] for an overview). However, none of these methods addresses cellwise contamination in the fuzzy clustering framework.
Figure 1: Data matrix with 250 observations, $10$ variables and $5\%$ of contaminated cells in black (left); corresponding rows in black trimmed via trimming approaches (right).
In this paper, we aim to fill this gap by combining the cellwise outlier detection with the advantages of fuzzy clustering. Indeed, after the detection of atypical values, the reliable cells may still provide useful information to assign observations to clusters in a fuzzy manner. We propose herein a cellwise fuzzy clustering method, called cellFCLUST, which intends to uncover a clustering structure that accommodates both hard and soft assignments on the one hand, and to robustly estimate the cluster parameters in the presence of cellwise contamination on the other. The levels of hard assignments are tuned by managing the scale of the data in a user-controlled manner. Along with its fuzzy nature, this is a key feature that makes cellFCLUST unique within the clustering framework with cellwise outliers.
The proposal is formulated via a maximum likelihood approach, assuming a Gaussian distribution for the clusters and allowing each to have a distinct shape. Therefore, the dependence relationships among variables may vary across clusters, and they are used for detecting and “correcting” the outlying cells. Specifically, by setting a proportion of unreliable cells for the variables, cellFCLUST identifies and imputes them using the information contained in the remaining reliable cells of each observation, making the use of an Expectation-Maximization (EM, [25]) type algorithm suitable for its estimation. As a result of the imputation, all observations are assigned to clusters with a certain degree – unlike trimming approaches, which discard observations, or noise clustering methods, where outliers are not grouped with regular data. Additionally, casewise methodologies, such as F-TCLUST, systematically trim observations in the tails of the clusters’ distributions, leading the estimators of the scatter matrices to be biased downward due to the removal of part of the clusters’ variability. For this reason, casewise estimators – even in the single-population framework – require consistency factor corrections to prevent this issue and, for instance, achieve Fisher consistency under normality. In contrast, cellFCLUST, as well as cellMCD for single populations, does not suffer from this problem, and correction factors for the scatter matrices are no longer required, since the application of the EM algorithm naturally corrects it.
The potential of the proposed methodology is demonstrated through a simulation study, in which we provide evidence of its strengths compared to other existing fuzzy clustering methods both in the non-robust and robust frameworks. Additionally, cellFCLUST is applied to two real data sets from different fields. In the first application, we detect groups of individuals with similar circumference measurements, which support the classification of subjects according to different body fat levels and, therefore, the identification of individuals at risk of obesity. The second application concerns regional data for studying well-being in the OECD countries.
@@
The rest of the paper is organized as follows. Section 2 introduces cellFCLUST, its parameter estimation, and the corresponding algorithm. A simulation study is implemented to evaluate the performance of cellFCLUST in comparison with alternative methods for fuzzy clustering with and without outlier detection, and the results are reported in Section 3. In Section 4, we provide guidance on the tuning parameter selection through examples that illustrate their impact on cluster recovery and outlier detection. Two real data applications are presented in Section 5. Finally, Section 6 concludes the paper with a discussion on the proposed methodology, as well as potential directions for future developments in the cellwise clustering literature.
@@ 2 Cellwise fuzzy clustering
Let $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\prime}$ be an $(n\times J)$ data matrix, where $\mathbf{x}_{i}=(x_{i1},\ldots,x_{iJ})$ records the values of the $J$ variables for the $i$th observation. Some cells $x_{ij}$ can be missing or outlying, with the latter resulting from a contaminating mechanism affecting single entries of $\mathbf{X}$. We generally refer to both types of values as unreliable cells. We denote by $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{n}]^{\prime}$ the $(n\times J)$ cellwise indicator matrix, where $w_{ij}=1$ if the cell is reliable, i.e., neither missing nor outlying, and $w_{ij}=0$ otherwise. Differently from the outlying values, the $(i,j)$-positions of the missing values in a data matrix are known, and therefore the corresponding zeros into $\mathbf{W}$ can be set a priori. According to $\mathbf{w}_{i}$, we partition $\mathbf{x}_{i}$ into $\mathbf{x}_{i[\mathbf{w}_{i}]}$ and $\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}$, where $\mathbf{w}_{i}^{c}=\mathbf{1}_{J}-\mathbf{w}_{i}$ with $\mathbf{1}_{J}$ being the unitary vector. These two sub-vectors represent the reliable and unreliable cells for the $i$th observation, respectively.
Our goal is to group the $n$ observations into $K$ clusters using a fuzzy approach, while handling contaminated data - and missing values, if present - and imputing them based on reliable information per observation and cluster-specific relationships among variables. The objective function of the fuzzy clustering problem with cellwise outlier detection, called cellwise Fuzzy Clustering (cellFCLUST), to maximize is
@@
where $\mathbf{U}$ is an $(n\times K)$ matrix that tracks the membership of each observation to a cluster, with entries $u_{ik}\in[0,1]$ and such that $\sum_{k=1}^{K}u_{ik}=1$ for $i=1,\ldots,n$; $m$ is the fuzzifier tuning parameter; $\phi_{J}(\cdot)$ is the probability density function of a $J$-variate normal distribution; and $J[\mathbf{w}_{i}]$ is the number of reliable cells for the $i$th observation. Both $\boldsymbol{m}_{1},\ldots,\boldsymbol{m}_{K}$, which are vectors in $\mathbb{R}^{J}$, and $\boldsymbol{\mathbf{S}}_{1},\ldots,\boldsymbol{\mathbf{S}}_{K}$, which are positive definite matrices of order $J$ holding the dependence relationships among variables within clusters, are restricted in (1) to the subvectors and the submatrices, respectively, corresponding to the reliable cells for the $i$th observation. The fuzzifier parameter $m$ must be greater than $1$ to obtain strictly fuzzy memberships, since $m=1$ results in crisp 010-1 assignments for every observation, even though $u_{ik}\in[0,1]$, as shown in [10]. The maximization of (1) is subject to the following constraints
@@
where $h$ and $c\geq 1$ are fixed constant, and $\lambda_{1}(\boldsymbol{\mathbf{S}}_{k}),\ldots,\lambda_{J}(\boldsymbol{\mathbf{S}}_{k})$ are the $J$ eigenvalues of $\boldsymbol{\mathbf{S}}_{k}$. Constraint (2) imposes that $h=\left\lceil(1-\alpha)n\right\rceil$ cells per variable are reliable. Here, $\alpha$ represents the proportion of flagged cells – for simplicity, we refer to this as “contamination level” throughout the paper – and it should be at most $0.25$ to guarantee that pairs of variables overlap for some observations, allowing their covariances to be computed. The eigenvalue-ratio constraint in (3) prevents unboundedness and spurious solutions that may arise during the maximization of the objective function. As the constant $c$ - which constraints the ratio between the largest and the smallest eigenvalues of the covariance matrices across clusters - decreases, the cluster configurations become increasingly similar, reaching spherical and equally scattered clusters when $c=1$.
It is worth noting that (1) may favor clusters with comparable values of $\sum_{i=1}^{n}u_{ik}^{m}$, which are quantities corresponding exactly to cluster sizes in a hard clustering approach ($m=1$). To avoid this effect and allow the sizes to vary across clusters, we may consider the maximization of the following objective function
@@
where $p_{k}\in(0,1)$, such that $\sum_{k=1}^{K}p_{k}=1$, are weights to estimate. In the remainder of the paper, we regard unbalanced clusters and therefore focus on the maximization of (4). Maximizing (4) rather than (1) is particularly relevant when the clusters have different sizes.
@@ 2.1 An EM-type algorithm for cellFCLUST
We aim to solve the maximization problem of the objective function in (4) via an EM-type algorithm similar to that used for handling missing data in the model-based clustering framework [26], with an additional step for cellwise outlier detection. Specifically, the algorithm is composed of four alternating steps. Its key feature lies in Step 1, where nhn-h cells per variable considered as potentially contaminated are flagged, and the corresponding positions in $\mathbf{W}$ are set to zero. The membership values in $\mathbf{U}$ are computed in Step 2. Following the EM rationale and treating the contaminated cells as missing, the parameters of the distribution for the missing part, necessary for the overall parameter estimation, are obtained in Step 3, resulting in the imputation of the contaminated cells, rather than their discard. Finally, Step 4 estimates the vectors $\boldsymbol{m}_{k}$ and the matrices $\boldsymbol{\mathbf{S}}_{k},k=1,\ldots,K$, subject to constraint (3). A detailed description of the algorithm is provided below.
@@
    Initial solutions for the parameters are obtained using several applications of TCLUST [27]. Specifically, TCLUST is applied to each variable and pairs of variables to initialize $\mathbf{W}$, and then on random subsets of variables to achieve feasible initializations for $p_{k},\boldsymbol{m}_{k}$, and $\boldsymbol{\mathbf{S}}_{k},k=1,\ldots,K$ (see [22] for more details). Accordingly, an initial solution for $\mathbf{U}$ is computed as described in Step 2.
@@
    Given the current parameters and the actual configuration of $\mathbf{W}$, we update the latter column-by-column. Let consider the contribution of the $i$th observation to the objective function in (4), i.e.
@@
    where $\hat{x}_{ij(k)}$ and $C_{ij(k)}$ are the expectation and the variance, respectively, of the cell corresponding to the $j$th variable in the $i$th observation, conditional on the other reliable cells for the $i$th observation and considering the parameters of the $k$th cluster. By sorting the $\Delta_{ij}$ values in an increasing order, i.e., $\Delta_{(1)j}\leq\Delta_{(2)j}\leq\cdots\leq\Delta_{(n)j}$, we flag as unreliable the observations with indexes $\{i:\Delta_{ij}<\Delta_{(n-h)j}\}$, where $h=\left\lceil(1-\alpha)n\right\rceil$, and set the corresponding elements of the $j$th column of $\mathbf{W}$ to zero. Equivalently, the observations with $\{i:\Delta_{ij}\geq\Delta_{(n-h)j}\}$ are considered reliable, which is denoted by ones into the $j$th columns of $\mathbf{W}$.
    If missing values occur for the $j$th variable, the corresponding cells of $\mathbf{W}$ are set to zero. Consequently, the $\Delta_{ij}$ values are computed only on the observed cells. This approach may result in varying proportions of zeros across variables, depending on the extent of missingness they contain. However, using $\alpha$ to account for both the contaminated and missing cells – and therefore grounding $h$ on $n$, independently of the number of missing values – could potentially lead to an incorrect identification of the truly contaminated cells. In general, we recommend excluding variables with a high proportion of missing data, particularly when moderate contamination is expected.
@@
    $\begin{cases}u_{ik}=I\Big\{f_{ik}=\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}\Big\}&\text{if}\,\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}\geq 1\\ u_{ik}=\Bigg(\sum_{k^{\prime}=1}^{K}\Bigg(\dfrac{\log(f_{ik})}{\log(f_{ik^{\prime}})}\Bigg)^{\frac{1}{m-1}}\Bigg)^{-1}&\text{if}\,\underset{k^{\prime}=1,\ldots,K}{\max}f_{ik^{\prime}}<1\end{cases},$ (7)
    where $f_{ik}=p_{k}\phi_{J[\mathbf{w}_{i}]}(\mathbf{x}_{i[\mathbf{w}_{i}]};\boldsymbol{m}_{k[\mathbf{w}_{i}]},\boldsymbol{\mathbf{S}}_{k[\mathbf{w}_{i},\mathbf{w}_{i}]})$ and $I\{\cdot\}$ represents the indicator function. In the first case, the $i$th observation is fully assigned to the $k$th cluster, while in the second case, the assignment is fuzzy. This is a desirable property of a fuzzy clustering approach since not all observations necessarily require a soft assignment. Some of them, especially those in the “core” of the clusters, can be unequivocally assigned.
@@
    Considering the MAR mechanism and the distinct parameter assumption [21], and thanks to the Gaussian distribution properties, we can decompose the probability density function of the complete data $\phi_{J}(\mathbf{x}_{i};\boldsymbol{m}_{k},$ $\boldsymbol{\mathbf{S}}_{k})$ into the product $\phi_{J[\mathbf{w}_{i}]}(\mathbf{x}_{i[\mathbf{w}_{i}]};\boldsymbol{m}_{k[\mathbf{w}_{i}]},$ $\boldsymbol{\mathbf{S}}_{k[\mathbf{w}_{i},\mathbf{w}_{i}]})\times\phi_{J[\mathbf{w}_{i}^{c}]}(\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}|\mathbf{x}_{i[\mathbf{w}_{i}]};$ $\boldsymbol{m}_{k[\mathbf{w}_{i}^{c}|\mathbf{w}_{i}]},\boldsymbol{\mathbf{S}}_{k[\mathbf{w}_{i}^{c},\mathbf{w}_{i}^{c}|\mathbf{w}_{i}]})$. Given the cellwise indicator matrix and the current parameters, we compute the parameters of the distribution of $\mathbf{x}_{i[\mathbf{w}_{i}^{c}]}$ conditional on $\mathbf{x}_{i[\mathbf{w}_{i}]}$ as
@@
    If the update of the covariance matrices in (12) does not satisfy the eigenvalue-ratio constraint in (3), we apply the efficient eigenvalue-truncation procedure proposed by Fritz et al. [28]. This ensures that the constraint in (3) is met.